# Emergent SU(3) topological system in a trimer SSH model

### v1

@@ Abstract
We consider a trimer Su-Schrieffer-Heeger (SSH) tight-binding Hamiltonian keeping up to next-nearest-neighbor (NNN) hopping terms and on-site potential energy. The Bloch Hamiltonian can be expressed in terms of all the eight generators (i.e. Gell-Mann matrices) of the SU(3) group. We provide exact analytical expressions of three dispersive energy bands and the corresponding eigenstates for any choices of the system parameters. The system lacks full chiral symmetry since the energy spectrum is not symmetric around zero, except at isolated Bloch wavevectors. We explore parity, time reversal, and certain special chiral symmetries for various system parameters. We discuss the bulk-boundary correspondence by numerically computing the Zak phase for all the bands and the boundary modes in the open boundary condition. There are three different kinds of topological phase transitions which are classified based on the gap closing points in the Brillouin zone (BZ) while tuning the nearest-neighbor (NN) and NNN hopping terms. We find that quantized changes (in units of $\pi$) in two out of three Zak phases characterize these topological phase transitions. We propose another bulk topological invariant, namely the sub-lattice winding number, which also characterizes the topological phase transitions changing from $\nu^{\alpha}=0\leftrightarrow 2$ and $\nu^{\alpha}=0\leftrightarrow 1\leftrightarrow 2$ ($\alpha$: sub-lattice index). The sub-lattice winding number provides a relatively simple analytical understanding of topological phases and may help in characterizing topological phases of systems without chiral symmetry.
@@ I Introduction
The geometrical and topological aspects of the Bloch bands in solid state systems have attracted intense research activities [1, 2, 3, 4, 5, 6]. The symmetry-protected topological (SPT) phases of matter can be understood in terms of the interplay between symmetry and topology [7, 8]. For non-interacting systems, the SPT phases are classified in the 10-fold periodic table according to the system’s spatial dimension and internal symmetries which include time reversal, particle-hole, and chiral symmetry [9]. The bulk-boundary correspondence is the underlying principle of topological phases which relates the non-trivial bulk topological invariant to the number of topological boundary states in the open boundary condition. Most of the studies are carried out in two-dimensional systems (2D) [10, 11], however the studies have been extended to one-dimensional (1D) [12, 13] and three-dimensional (3D) systems [14] as well.
@@
The simplest model describing the topological features is the dimer Su-Schrieffer-Heeger (SSH) model [15] consists of a 1D tight-binding model with alternating hopping parameters. Further, topological phases of the extended SSH model including the next-nearest-neighbour (NNN) hopping has been studied extensively [16]. The topological phase of a finite system can be characterized by the presence of zero energy boundary modes which are exponentially localized at the edges. The conventional bulk-boundary correspondence using the Zak phase [17], similar to the Berry phase [18], has been used as a topological number to classify various inversion-symmetric 1D topological systems where it is quantized to $\pi$ or 0 (mod $2\pi$) [19]. The boundary modes and the Zak phase in various artificial 1D systems such as optical lattices [20], acoustic crystals [21], photonic lattices [22], photonic crystals [23, 24] and artificial spin chains in ultracold atomic system [25] have been studied elaborately. However, there is an ambiguity to use the Zak phase as a topological number, since it depends on the choice of the origin in the real-space and how the unit cell is chosen [20, 26]. For systems in odd spatial dimensions with chiral and/or inversion symmetry, three bulk-topological invariants namely the winding number, and inversion symmetry indicator are proposed [27].
@@
The momentum space Hamiltonian of 1D, 2D and 3D topological systems having two bands is represented by irreducible SU(2) generators. However, there is a growing interest in three-band in 2D and 3D topological systems, whose momentum space Hamiltonian is represented by an irreducible SU(3) generators [28, 29] given in Appendix A. For example, Lieb [30, 31, 32] and $\alpha$-$\mathcal{T}_{3}$ lattices [33, 34] in 2D as well as triple-component fermions [35, 36, 37, 38] in 3D exhibits three bands, two dispersive bands, and one flat band. All three bands touch at a given ${\bm{k}}$ point in the Brillouin zone (BZ), leading to a three-fold degeneracy at the band touching point. In Ref. [39], it was proposed that in a square optical lattice with spin-orbit coupling subjected to a spatially homogeneous SU(3) synthetic gauge field can exhibit nontrivial topology. The possibility of designing a 2D lattice model having SU(3) symmetry for spinless fermion has been discussed in Ref. [40]. In a recent study [41], a two-dimensional electron gas with Rashba spin-orbit coupling has been generalized to an SU(3) system, by considering trigonal crystal field effect which leads to a partial splitting of the energy levels at the time-reversal invariant momenta. But there are very few studies [42, 43] on three-band topological systems in 1D.
@@
In this work, we consider a trimer SSH model by considering three sites in a given unit cell. Keeping on-site energy and hopping term up to NNN sites, the Bloch Hamiltonian is expressed in terms of all the eight generators of the SU(3) group. We obtain exact analytical expressions of three dispersive energy bands and the associated eigenstates for arbitrary choices of the system parameters. We explore various symmetry properties of the trimer SSH model for various system parameters. Since the system does not possess chiral symmetry, we establish the bulk-boundary condition by numerically computing the Zak phases ($\phi_{\lambda}$ with $\lambda=1,2,3$) for all the bands along with the boundary modes in the open boundary condition. Depending on the gap closing points in the BZ, we observe three different kinds of topological phase transitions while tuning NN and NNN hopping terms. We find that quantized changes in two out of three Zak phases are necessary to characterize these topological phase transitions, namely $\phi_{1}+\phi_{3}$ changing from $\phi_{1}+\phi_{3}=0\leftrightarrow 2\pi$ as well as $\phi_{1}+\phi_{3}=0\leftrightarrow\pi\leftrightarrow 2\pi$ characterize these topological phase transitions. Moreover, we propose another topological invariant, namely the sub-lattice winding number which also characterizes topological phase transitions similar to the Zak phases. We derive exact analytical expressions of the sub-lattice winding number which changes from $\nu^{\alpha}=0\leftrightarrow 2$ as well as $\nu^{\alpha}=0\leftrightarrow 1\leftrightarrow 2$ across the topological phase transitions.
@@
This paper is organized as follows. In Sec. II, we describe the trimer SSH model and obtain the periodic Bloch Hamiltonian, dispersive energy bands along with the corresponding eigenstates and various symmetries of the system. In Sec. III, we discuss Zak phases, sub-lattice winding number, and the boundary modes in the open boundary condition. We summarize our results in Sec. IV. In Appendix A, we provide an explicit form of the Gell-Mann matrices. We obtain canonical Bloch Hamiltonian and show the equivalence with the periodic Bloch Hamiltonian in Appendix B. The phase diagram characterizing topological phases in different parameter regimes is given in Appendix C. The sub-lattice winding number for the dimer SSH model is discussed in Appendix D.
@@ II SU(3) Hamiltonian and its various properties
Here the operators $a_{l}^{\dagger}$, $b_{l}^{\dagger}$ and $c_{l}^{\dagger}$ create a spinless fermion at the sub-lattice sites $A,B$ and $C$ in the $l$-th unit cell of unit length and $h.c.$ stands for the hermitian conjugate. As opposed to the dimer SSH model, the NNN hopping occurs between two different sites in the same cell as well as in the next cell. In the absence of NNN hopping amplitudes and on-site energy, the above Hamiltonian $H$ reduces to that of in Ref. [42, 43].
@@
Here ${\bm{\Lambda}}$ is a vector composed of the Gell-Mann matrices and ${\bm{d}(k)}$ is an eight-dimensional real vector arising due to various hopping terms up to NNN sites in 1D, whose components are $d_{1}=u+u_{1}\cos(k),d_{2}=u_{1}\sin(k),d_{3}=\Delta/2,d_{4}=w_{1}+w\cos(k),d_{5}=w\sin(k),d_{6}=v+v_{1}\cos(k),d_{7}=v_{1}\sin(k)$ and $d_{8}=\sqrt{3}\Delta/2$. The length of the eight-dimensional vector ${\bm{d}(k)}$ is $L(k)=\sqrt{\Delta^{2}+|U|^{2}+|V|^{2}+|W|^{2}}$. In the absence of the NNN hopping, $L$ becomes $k$-independent, which is in sharp contrast to the dimer SSH model.
@@ II.1 Energy bands and their properties
The Hamiltonian, $\mathcal{H}(k)$ yield the following depressed cubic equation for eigenvalues $E$: $E^{3}+pE+q=0$ with
$p$ $=$ $-[(\Delta^{2}+u^{2}+v^{2}+w^{2}+u_{1}^{2}+v_{1}^{2}+w_{1}^{2})+2(uu_{1}+vv_{1}+ww_{1})\cos(k)]=-L^{2}(k)$ (9)
$q$ $=$ $-[\Delta(u^{2}-v^{2}+u_{1}^{2}-v_{1}^{2})+2(u_{1}vw+uv_{1}w+uvw_{1})+2u_{1}v_{1}w_{1}\cos(2k)$ (10)
$+$ $2(\Delta(uu_{1}-vv_{1})+uvw+u_{1}v_{1}w+u_{1}vw_{1}+uv_{1}w_{1})\cos(k)].$
It should be noted that $p$ is always negative since $p=-L^{2}(k)$. The three energy bands are
@@
with $\lambda=1,2,3$ and $E_{3}\geq E_{2}\geq E_{1}$. It is important to note here that the middle band $E_{2}(k)$ vanishes at two $k$ points whenever $q=0$. Therefore the two locations of the zeros of $E_{2}(k)$ can be found from Eq. (10) for the given system parameters. For example, $E_{2}(k=\pm\pi/2)=0$ for the following two simple cases: (i) $\Delta\neq 0$, $u=v$ and $u_{1}=v_{1}=w_{1}=0$ and (ii) for any values of $u,v,w$ with $\Delta=u_{1}=v_{1}=w_{1}=0$. Figures 2 and 3 display the band dispersion for the various system parameters.
@@
for any given $k$. In three-dimensional Euclidean space with the axes $E_{1},E_{2},E_{3}$, Eq. (13) describes a plane passing through the origin, for which the normal line has direction cosines: $(1/\sqrt{3})(1,1,1)$. Similarly, Eq. (14) is a sphere of radius $\sqrt{2}L(k)$ centered at the origin.
@@ II.2 Symmetry properties of the Hamiltonian $\mathcal{H}(k)$
The concept of point chiral symmetry has been introduced for $\Delta=u_{1}=v_{1}=w_{1}=0$ case in Ref. [43, 44, 45]. Mathematically, the point chiral symmetry is stated as $\Gamma_{p}\mathcal{H}(k)\Gamma^{\dagger}=-\mathcal{H}(k+\pi)$ with the point chiral operator $\Gamma_{p}=\text{diag}(1,-1,1)$. It immediately follows $E_{1}(\pi-k)=-E_{3}(k)$ and $E_{2}(\pi-k)=-E_{2}(k)$. We extend the concept of point chiral symmetry for $\Delta\neq 0,u_{1}=v_{1}=w_{1}=0$ along with $u=v$ case. We find that the point chiral symmetry is preserved even in the presence of $\Delta$ with the condition $u=v$. The eigenstates of the shifted particle-hole symmetry are connected as $\Gamma_{p}\psi_{1}(\pi-k)=\psi_{3}(k)$ and $\Gamma_{p}\psi_{2}(\pi-k)=\psi_{2}(k)$.
The NNN hopping terms break both the chiral symmetries due to the presence of the Bloch phases ($\sim e^{\pm ik}$) arising from the inter-cell and intra-cell NNN hopping terms.
III Topological phases and bulk-boundary correspondence
In 1D, no non-trivial topological phase is possible for systems having no symmetry. For systems in odd spatial dimensions with chiral and/or inversion symmetry, three bulk-topological invariants namely the winding number ($\nu$), Zak phase ($\phi_{\lambda}$), and inversion symmetry indicator ($\chi$) are proposed [27]. These topological invariants can take non-trivial values depending on the symmetries of the system and correspond to boundary modes in the open boundary condition. The winding number is defined as follows [46]

### v2

# Bulk-boundary correspondence in extended trimer Su-Schrieffer-Heeger model

@@ Abstract
We consider an extended trimer Su-Schrieffer-Heeger (SSH) tight-binding Hamiltonian keeping up to next-nearest-neighbor (NNN) hopping terms and on-site potential energy. The Bloch Hamiltonian can be expressed in terms of all the eight generators (i.e. Gell-Mann matrices) of the SU(3) group. We provide exact analytical expressions of three dispersive energy bands and the corresponding eigenstates for any choices of the system parameters. The system lacks full chiral symmetry since the energy spectrum is not symmetric around zero, except at isolated Bloch wavevectors. We explore parity, time reversal, and certain special chiral symmetries for various system parameters. We discuss the bulk-boundary correspondence by numerically computing the Zak phase for all the bands and the boundary modes in the open boundary condition. There are three different kinds of topological phase transitions, which are classified based on the gap closing points in the Brillouin zone (BZ) while tuning the nearest-neighbor (NN) and NNN hopping terms. We find that quantized changes (in units of $\pi$) in two out of three Zak phases characterize these topological phase transitions. We propose another bulk topological invariant, namely the sub-lattice winding number, which also characterizes the topological phase transitions changing from $\nu^{\alpha}=0\leftrightarrow 2$ and $\nu^{\alpha}=0\leftrightarrow 1\leftrightarrow 2$ ($\alpha$: sub-lattice index). The sub-lattice winding number not only provides a relatively simple analytical understanding of topological phases but also successfully establishes bulk-boundary correspondence in the absence of inversion symmetry, which may help in characterizing the bulk-boundary correspondence of systems without chiral and inversion symmetry.
@@ I Introduction
The geometrical and topological aspects of the Bloch bands in solid state systems have attracted intense research activities [1, 2, 3, 4, 5, 6]. The symmetry-protected topological (SPT) phases of matter can be understood in terms of the interplay between symmetry and topology [7, 8]. For non-interacting systems, the SPT phases are classified in the 10-fold periodic table according to the system’s spatial dimension and internal symmetries, which include time reversal, particle-hole, and chiral symmetries [9, 10]. The bulk-boundary correspondence is the underlying principle of topological phases which relates the non-trivial bulk topological invariant to the number of topological boundary modes in the open boundary condition. Most of the studies are carried out in two-dimensional systems (2D) [11, 12], however the studies have been extended to one-dimensional (1D) [13, 14] and three-dimensional (3D) systems [15] as well.
@@
The simplest model describing the topological features is the dimer Su-Schrieffer-Heeger (SSH) model [16] consists of a 1D tight-binding model with alternating hopping parameters. Further, topological phases of the extended SSH model [17], periodically driven 1D optical lattice [18] and quasi-1D lattice [19, 20] including the next-nearest-neighbour (NNN) hopping have been studied extensively. The topological phase of a finite system can be characterized by the presence of zero energy boundary modes which are exponentially localized at the edges. The conventional bulk-boundary correspondence using the Zak phase [21], similar to the Berry phase [22], has been used as a topological number to classify various inversion-symmetric 1D topological systems where it is quantized to $\pi$ or 0 (mod $2\pi$) [23]. The boundary modes and the Zak phase in various artificial 1D systems such as optical lattices [24], acoustic crystals [25], photonic lattices [26], photonic crystals [27, 28] and artificial spin chains in ultracold atomic system [29] have been studied elaborately. However, there is an ambiguity to use the Zak phase as a topological number, since it depends on the choice of the origin in the real-space and how the unit cell is chosen [24, 30]. For systems in odd spatial dimensions with chiral and/or inversion symmetry, three bulk topological invariants namely the winding number, the Zak phase, and inversion symmetry indicator are proposed [31].
@@
The momentum space Hamiltonian of 1D, 2D and 3D topological systems having two bands is represented by irreducible SU(2) generators. However, there is a growing interest in three-band in 2D and 3D topological systems, whose momentum space Hamiltonian is represented by an irreducible SU(3) generators [32, 33] given in Appendix A. For example, Lieb [35, 36, 37] and $\alpha$-$\mathcal{T}_{3}$ lattices [38, 39] in 2D as well as triple-component fermions [40, 41, 42, 43] in 3D exhibits three bands, two dispersive bands, and one flat band. All three bands touch at a given ${\bm{k}}$ point in the Brillouin zone (BZ), leading to a three-fold degeneracy at the band touching point. In Ref. [34], the geometric phase structure of SU(3) fermions in graphene-like models has also been discussed. In Ref. [44], it was proposed that in a square optical lattice with spin-orbit coupling subjected to a spatially homogeneous SU(3) synthetic gauge field can exhibit nontrivial topology. The possibility of designing a 2D lattice model having SU(3) symmetry for spinless fermion has been discussed in Ref. [45]. In a recent study [46], a two-dimensional electron gas with Rashba spin-orbit coupling has been generalized to an SU(3) system, by considering trigonal crystal field effect which leads to a partial splitting of the energy levels at the time-reversal invariant momenta. But there are very few studies [19, 20, 47, 48, 49] on multi-band topological systems in 1D.
@@
In this work, we consider a trimer SSH model by considering three sites in a given unit cell. Keeping on-site energy and hopping term up to NNN sites, the Bloch Hamiltonian is expressed in terms of all the eight generators of the SU(3) group. We obtain exact analytical expressions of three dispersive energy bands and the associated eigenstates for arbitrary choices of the system parameters. We explore various symmetry properties of the trimer SSH model for various system parameters. Since the system does not possess chiral symmetry, we establish the bulk-boundary condition by numerically computing the Zak phases ($\phi_{\lambda}$ with $\lambda=1,2,3$) for all the bands along with the boundary modes in the open boundary condition. Depending on the gap closing points in the BZ, we observe three different kinds of topological phase transitions while tuning NN and NNN hopping terms. We find that quantized changes in two out of three Zak phases are necessary to characterize these topological phase transitions, namely $\phi_{1}+\phi_{3}$ changing from $\phi_{1}+\phi_{3}=0\leftrightarrow 2\pi$ as well as $\phi_{1}+\phi_{3}=0\leftrightarrow\pi\leftrightarrow 2\pi$ characterize these topological phase transitions. Moreover, we propose another topological invariant, namely the sub-lattice winding number, which also characterizes topological phase transitions similar to the Zak phases. We derive exact analytical expressions of the sub-lattice winding number, which changes from $\nu^{\alpha}=0\leftrightarrow 2$ as well as $\nu^{\alpha}=0\leftrightarrow 1\leftrightarrow 2$ across the topological phase transitions. We find that the sub-lattice winding number provides a relatively simple analytical understanding of topological phases and also successfully establishes the bulk-boundary correspondence without inversion asymmetric systems, which may help characterize the bulk-boundary correspondence of systems without chiral and inversion symmetry.
@@
This paper is organized as follows. In Sec. II, we describe the trimer SSH model and obtain the periodic Bloch Hamiltonian dispersive energy bands along with the corresponding eigenstates and various symmetries of the system. In Sec. III, we discuss the bulk topological invariants Zak phases and sub-lattice winding numbers. In Sec. IV and V, we discuss the bulk-boundary correspondences in the absence and presence of next-nearest neighbor hopping, respectively. We summarize our results in Sec. VI. In Appendix A, we provide an explicit form of the Gell-Mann matrices. We obtain canonical Bloch Hamiltonian and show the equivalence with the periodic Bloch Hamiltonian in Appendix B. The phase diagram characterizing topological phases in different parameter regimes is given in Appendix C. The sub-lattice winding number for the dimer SSH model is discussed in Appendix D. The real space profiles of the edge states are discussed in Appendix E.
@@ II SU(3) Hamiltonian and its various properties
Here the operators $a_{l}^{\dagger}(a_{l})$, $b_{l}^{\dagger}(b_{l})$ and $c_{l}^{\dagger}(c_{l})$ create (destroy) a spinless fermion at the sub-lattice sites $A,B$ and $C$ in the $l$-th unit cell of unit length and $h.c.$ stands for the hermitian conjugate. As opposed to the dimer SSH model, the NNN hopping occurs between two different sites in the same cell as well as in the next cell. In the absence of NNN hopping amplitudes and on-site energy, the above Hamiltonian $H$ reduces to that of in Ref. [47, 48].
@@
Here ${\bm{\Lambda}}$ is a vector composed of the Gell-Mann matrices and ${\bm{d}(k)}$ is an eight-dimensional real vector arising due to various hopping terms up to NNN sites in 1D, whose components are $d_{1}=u+u_{1}\cos(k),d_{2}=u_{1}\sin(k),d_{3}=\Delta/2,d_{4}=w_{1}+w\cos(k),d_{5}=w\sin(k),d_{6}=v+v_{1}\cos(k),d_{7}=v_{1}\sin(k)$ and $d_{8}=\sqrt{3}\Delta/2$. The length of the eight-dimensional vector ${\bm{d}(k)}$ is $L(k)=\sqrt{\Delta^{2}+|U|^{2}+|V|^{2}+|W|^{2}}$. In the absence of the NNN hopping, $L(k)$ becomes $k$-independent, which is in sharp contrast to the dimer SSH model.
@@ II.1 Energy bands and their properties
The Hamiltonian, $\mathcal{H}(k)$ yield the following depressed cubic equation for eigenvalues $E$: $E^{3}+pE+q=0$ with $p=-L^{2}(k)$ and $q=-[(|U|^{2}-|V|^{2})\Delta+UVW^{*}+U^{*}V^{*}W]$. The three energy bands are
@@
with $\lambda=1,2,3$ and $E_{3}\geq E_{2}\geq E_{1}$. It is important to note here that the middle band $E_{2}(k)$ vanishes at two $k$ points whenever $q=0$, which will be useful in discussing the symmetries of the system. Therefore the two locations of the zeros of $E_{2}(k)$ can be found from the expression of $q$ for the given system parameters. For example, $E_{2}(k=\pm\pi/2)=0$ for the following two simple cases: (i) $\Delta\neq 0$, $u=v$ and $u_{1}=v_{1}=w_{1}=0$ and (ii) for any values of $u,v,w$ with $\Delta=u_{1}=v_{1}=w_{1}=0$. Figures 2 and 3 display the band dispersion for the various system parameters.
@@
for any given $k$. In three-dimensional Euclidean space with the axes $E_{1},E_{2},E_{3}$, Eq. (11) describes a plane passing through the origin, for which the normal line has direction cosines: $(1/\sqrt{3})(1,1,1)$. Similarly, Eq. (12) is a sphere of radius $\sqrt{2}L(k)$ centered at the origin. These properties will be useful in discussing the symmetries of the system.
@@ II.2 Symmetry properties of the Hamiltonian $\mathcal{H}(k)$
II.2.1 $\Delta=0$
For $u=v$ and in the presence of NNN hopping $w_{1}$ and $u_{1}=v_{1}$, the space-inversion symmetry is preserved. This is quite different from the dimer SSH model, where the space-inversion symmetry is broken as soon as we consider different NNN hopping terms for different sub-lattices since they involve hopping between two successive equivalent sites result in a diagonal element in the Bloch Hamiltonian. In the case of the trimer SSH model, hopping between two successive equivalent sites are only possible when next to NNN hopping is considered. The space-inversion symmetry will be broken in the trimer SSH model if we consider the next NNN hopping terms. Moreover, the inversion symmetry is broken for $u\neq v$ and or $u_{1}\neq v_{1}$ for any value of $w_{1}$.
The system lacks full chiral symmetry since the energy bands are not symmetric around zero for all values of $k$. The concept of point chiral symmetry for $u\neq v$ (inversion asymmetric) has been introduced for $\Delta=u_{1}=v_{1}=w_{1}=0$ case in Ref. [48, 50, 51]. Mathematically, the point chiral symmetry is stated as $\Gamma_{p}\mathcal{H}(k)\Gamma^{\dagger}=-\mathcal{H}(k+\pi)$ with the point chiral operator $\Gamma_{p}=\text{diag}(1,-1,1)$. It immediately follows $E_{1}(\pi-k)=-E_{3}(k)$ and $E_{2}(\pi-k)=-E_{2}(k)$. Moreover, we find that the point chiral symmetry is broken in the presence of NNN hopping ($u_{1},v_{1},w_{1}\neq 0$).
II.2.2 $\Delta\neq 0$
@@
We extend the concept of point chiral symmetry for $\Delta\neq 0,u_{1}=v_{1}=w_{1}=0$ along with $u=v$ case. We find that the point chiral symmetry is preserved even in the presence of $\Delta$ with the condition $u=v$. The eigenstates of the shifted particle-hole symmetry are connected as $\Gamma_{p}\psi_{1}(\pi-k)=\psi_{3}(k)$ and $\Gamma_{p}\psi_{2}(\pi-k)=\psi_{2}(k)$. It should be noted here that the Bloch phases ($\sim e^{\pm ik}$) arising from the inter-cell and intra-cell NNN hopping terms break the point chiral symmetry.
III Topological invariants
In 1D, no non-trivial topological phase is possible for systems having no symmetry. For systems in odd spatial dimensions with chiral and/or inversion symmetry, three bulk-topological invariants, namely the winding number ($\nu$), Zak phase ($\phi_{\lambda}$), and inversion symmetry indicator ($\chi$) are proposed [31]. These topological invariants can take non-trivial values depending on the symmetries of the system and correspond to boundary modes in the open boundary condition. The bulk winding number is defined as follows [52]